OpenStudy (anonymous):

If xy = 6 and x^2 + y^2 = 16, then what is the value of (x + y)?

OpenStudy (anonymous):

X+y=5

OpenStudy (anonymous):

please can you explain

OpenStudy (anonymous):

im trying to remember how to do this one sec

OpenStudy (anonymous):

\[x ^{2}+y ^{2}+2xy =16+6\times2\] \[\left( x +y \right)^{2}=28\] \[x +y =\pm \sqrt{28}\]

OpenStudy (anonymous):

dunno how you came up with that but no way that works

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

its obvious the values of x and y are 2 and 3 but i cannot for the life of me remember how to explain it

OpenStudy (anonymous):

so x+y=5

myininaya (myininaya):

\[(x+y)^2=x^2+2xy+y^2=x^2+y^2+2(xy)=16+2(6)=16+12=28\] so we have \[(x+y)^2=28 => x+y=\pm \sqrt{28}\]

myininaya (myininaya):

so i agree with lazy :)